We have the following information about a country’s economic performance in 3 consecutive years:

 GDP at current prices

GDP at previous year’s prices

 

year 0

20 000

18 000

 

 

year 1

30 000

25 000

 

 

year 2

45 000

40 000

 

 

 

Answer the following questions:

 

a)  From year 0 to year 2 the nominal GDP increased by %.

b)  Based on the GDP deflator the prices increased % from year 0 to year 2.

c)  The real GDP increase from year 0 to year 2 is % (rounded to one decimal).

d)  In year 0 the inflation rate (price increase relative to the previous year) was % (rounded to one decimal again).


1
Expert's answer
2021-04-02T07:25:00-0400

a)  Increase in nominal GDP (from year 0 to year 2) = (45000 - 20000) / 20000 = 125.0%


b)  GDP deflator (year 0) = (Nominal GDP / Real GDP) * 100 = (20000 / 18000) * 100 = 111.1

GDP deflator (year 2) = (45000 / 40000) * 100 = 112.5

Increase in prices (from year 0 to year 2) = 112.5 - 111.1 = 1.4%


c)  Increase in real GDP (from year 0 to year 2) = (40000 - 18000) / 18000 = 122.2%


d)  GDP deflator can be useful to determine inflation rate in year 0. GDP deflator in year 0 was 111.1 that means inflation rate was 11.1%. This means the price increase relative to the previous year was 11.1% in year 0.